After 10 s an object is displaced 40 m [W] due to a uniform acceleration of - 5 m/s squared. What is the object's final velocity?
1
Expert's answer
2011-11-03T12:23:08-0400
Let's find the initial velocity Vi of an object. We know that
Vi*t + (a*t^2)/2 = d
So, Vi = (d - (a*t^2)/2)/t = d/t + a*t/2 = 40/10 + 5*10/2 = 29 m/s.
Now we can find the object's final velocity Vf:
Vf = Vi + a*t = 29 + 5*10 = 79 m/s.